x+y+z=9

2x-3y+4z=13

3x+4y+5z=40


1
Expert's answer
2021-01-06T18:42:36-0500

x + y + z = 9

2x - 3y + 4z = 13

3x + 4y + 5z = 40   to find x, y, z we use Row echelon form

"\begin{bmatrix}
 1 & 1 & 1 && 9 \\
 2 & -3 & 4 && 13 \\ 
3 & 4 & 5 && 40 \\ 
\end{bmatrix}" = "\begin{vmatrix}
 
 R_2 = 2R_1 - R_2 \\ 
R_3 = 2R_3-3R_2 \\ 
\end{vmatrix}" =


"\begin{bmatrix}
 1 & 1 & 1 && 9 \\
 0 & 5 & -2 && 5 \\ 
0 & 17 & -2 && 41 \\ 
\end{bmatrix}" = "\begin{vmatrix}
R_3 = 5R_3-17R_2 \\ 
\end{vmatrix}" =


"\begin{bmatrix}
 1 & 1 & 1 && 9 \\
 0 & 5 & -2 && 5 \\ 
0 & 0 & 24 && 120 \\ 
\end{bmatrix}" = "\begin{bmatrix}
 1 & 1 & 1 && 9 \\
 0 & 1 & -0.4 && 1 \\ 
0 & 0 & 1 && 5 \\ 
\end{bmatrix}" => z = 5;


y - 0.4z = 1 => y - 0.4*5 = 1 => y =  3;

x + y + z = 9 => x + 3 + 5 = 9 => x = 1;


Answer: [1, 3, 5]